Suppose that the market demand for aluminum in a given country is given by the equation Qd =600-50P WHERE P is the price of aluminum expressed in dollar per kilogram and want to adopt price


1
Expert's answer
2021-06-28T01:19:03-0400

Solution


Given the following;


Qd= 600 - 50P and Qs = -400 + 50P


At equilibrium, Qd = Qs


600 - 50P = -400 + 50p


600 + 400 = 50P + 50P


1000 = 100P


"\\frac{1000}{100} = P \n\n\n\u200b"


P = 10


Equlibrium price = $10


Equilibrium Quantinty = "600 - (50 \times10) or -400 + (50 \times 10)"


= 600 - 500 or - 400 + 500 = 100units


After 10% increase, New Price = "10 + (10 \times0.1)= 11"


Effects on Quantity Demanded and Quantity supplied:


Qd = 600 -"(50 \times 11)" = 600 - 550 = 50units (decrease)


Qs = -400 + "(50 \times11)" = -400 + 550 = 150units (increase)